OpenStudy (anonymous):

Solving Radical Equations 4^ square root x-7=2

OpenStudy (anonymous):

\[4^{\sqrt{x}}-7=2\]

OpenStudy (anonymous):

\[4^{\sqrt{x}}=9\]

OpenStudy (anonymous):

the 4 is ^4

OpenStudy (anonymous):

that is what I don't know how solve

OpenStudy (midhun.madhu1987):



OpenStudy (anonymous):

\[^{4\sqrt{x-7=2}}\]

OpenStudy (midhun.madhu1987):

 now square both sides...